A cyclotron, designed to accelerate protons, has a magnetic field of 0.450 T magnitude in a region of radius 1.20 m. What are the values ​​of (a) the frequency of the cyclotron, (b) the maximum speed acquired by the protons, (c) the acceleration experienced by the protons and (d) the magnetic force exerted on the protons?
Answer #1

Use formulae carefully(a) w = 274 - IB m = 1.67x10-27 1.600 m for proton 9. x 100% С kg 274 1:60 x 10-19 X: 450 1.67* 10-27 f = 6.8617 MHZ (b) q Br